OpenStudy (anonymous):

Please help!!!!!! solve : 4x - 2x - 2 = -42.

OpenStudy (anonymous):

Any help is really appreciated!

OpenStudy (anonymous):

What you want to do for these types of questions is to isolate your variable 'x' on one side and the rest without your variable on the other.

OpenStudy (anonymous):

so like you simplify to 2x - 2 = -42 first right

OpenStudy (anonymous):

Yup. You can simplify first and then move the two on the other side by adding 2 to both sides.

OpenStudy (anonymous):

We add 2 to both sides because we want to balance the equation. Whatever we to do one side, we have to do the same with the other.

OpenStudy (anonymous):

so 2x - 2 + 2 = -42 + 2?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so then 2x = -40...

OpenStudy (anonymous):

Since we want to find what is x, what is the operation to get x and not 2x.

OpenStudy (anonymous):

Oh right you do -40 / 2 right?

OpenStudy (anonymous):

Yup. That's correct

OpenStudy (anonymous):

so then it would be x = -20

OpenStudy (anonymous):

That is the final answer

OpenStudy (anonymous):

awesome thank you so much!

OpenStudy (anonymous):

Cheers